OpenStudy (anonymous):

solve for a: v = v0 + at

OpenStudy (anonymous):

I'm guessing that v0 is one variable. v = v0 + at v0+at=v -v0 -v0 at = v - v0 divide both sides by t a = (v-V0)/t